### v1

# Imitation Learning with Limited Actions via Diffusion Planners and Deep Koopman Controllers

Abstract
Recent advances in diffusion-based robot policies have demonstrated significant potential in imitating multi-modal behaviors. However, these approaches typically require large quantities of demonstration data paired with corresponding robot action labels, creating a substantial data collection burden. In this work, we propose a plan-then-control framework aimed at improving the action-data efficiency of inverse dynamics controllers by leveraging observational demonstration data. Specifically, we adopt a Deep Koopman Operator framework to model the dynamical system and utilize observation-only trajectories to learn a latent action representation. This latent representation can then be effectively mapped to real high-dimensional continuous actions using a linear action decoder, requiring minimal action-labeled data. Through experiments on simulated robot manipulation tasks and a real robot experiment with multi-modal expert demonstrations, we demonstrate that our approach significantly enhances action-data efficiency and achieves high task success rates with limited action data.
I Introduction
Learning effective policies from demonstrations is a fundamental challenge in robotics and machine learning. Traditional imitation learning often relies on complete state-action trajectories, but obtaining action labels can be costly or impractical in many real-world scenarios. This limitation has spurred interest in imitation from observation, where the goal is to learn policies primarily from state observations without corresponding action data. Successfully addressing this challenge would enable scalable learning systems capable of leveraging abundant unlabeled observational data to perform complex tasks.
In this work, we adopt a plan-then-control scheme with Decision Diffusers[1] (Fig. 1) and propose Koopman Operator with Action Proxy (KOAP), a simple yet effective approach for learning an inverse dynamics controller with limited action data. Inspired by Deep Koopman Operators[2], the key insight underlying KOAP is to learn useful latent actions by regularizing the latent space with linear dynamics. By enforcing linear forward dynamics, the learned latent actions capture meaningful differences between lifted states, facilitating accurate action predictions even with limited action supervision. Applying a simple (linear) action decoder further minimizes overfitting risks when action-labeled data is scarce.
Figure 1: We adopt a plan-then-control scheme: a diffusion-based planner generates future states based on current and past states, with an inverse dynamics controller that generate action sequence to follow target trajectory. We propose KOAP, a method for leveraging action-free trajectories to improve controller learning with limited action data. KOAP exploits Deep Koopman Operators to lift the nonlinear target system into a linear latent space, which regularizes latent action learning. Real actions can be decoded through learning a simple (linear) action decoder by using action data.
Our approach differs from prior works such as ILPO[3] and LAPO[4], which focus primarily on discrete action policies and involve more complex modeling of latent actions and dynamics, often requiring information bottlenecks to retain essential information. In contrast, KOAP lifts the nonlinear system into a linear space of continuous latent states and latent actions, allowing us to naturally handle continuous action policies. By integrating decision diffusers for planning and focusing on learning an effective inverse model, we directly map latent actions to real actions without the need for complex architectures or quantization bottlenecks. To our knowledge, KOAP is the first work to combine decision diffusers with Koopman-inspired inverse model learning, providing a simple and scalable solution for imitation learning with limited action data.
We validate KOAP on a range of simulated robot manipulation tasks from the D3IL benchmark[5]. Our experiments demonstrate that KOAP significantly outperforms baseline methods, particularly when action-labeled data is limited. We observe that KOAP effectively learns informative latent actions, resulting in strong policy performance. Performance improves with additional observational data, and KOAP can even learn useful latent representations without any action labels. A case-study for scooping on a real-robot system shows our approach outperforms existing methods.
By addressing the challenge of learning from observations with limited action data, KOAP advances the field of imitation from observations. Our key contributions are:
  • We introduce KOAP, a simple, scalable method that combines decision diffusers for planning with effective inverse model learning, leveraging Koopman Operator theory to regularize latent actions through linear latent dynamics and linear action decoding.
  • We demonstrate that KOAP effectively learns informative latent actions, leading to strong policy performance even with limited action-labeled data.
  • We show that KOAP outperforms existing methods and variants on complex manipulation tasks.
This work enables scalable and efficient policy learning from observations, reducing dependence on costly action labeling and paving the way for more flexible imitation learning systems capable of handling continuous action spaces.
II Related Work
In this work, we focus on learning policies from demonstrations using primarily state observations rather than full state-action trajectories. This area is commonly known as imitation from observation or state-only imitation learning. It has seen significant growth in recent years, and we refer interested readers to recent surveys for comprehensive coverage (e.g.,[6, 7]). In brief, research in this area spans various approaches, including learning goals[8, 9, 10], reward functions[11], or trajectory waypoints[12] from state-only demonstrations. To enhance generalization across different visual settings and tasks, prior studies have focused on learning visual dynamics[13, 14, 15] or visual representations[16, 17]. Some approaches also incorporate auxiliary signals, such as rewards[18, 15] or interactions with the environment[19, 20, 21].
Our work is related to methods that learn inverse dynamics models. For example, VPT[22] trains an inverse dynamics model as an action labeler using labeled trajectories. VPT uses this model to assign pseudo-actions to action-free trajectories (videos), achieving human-level performance for Minecraft agents on specific tasks. However, unlike KOAP, VPT does not use a forward dynamics model and risks overfitting the action labeler when few action trajectories are available.
The most relevant works to ours are latent action learning methods ILPO[3] and LAPO[4]. Unlike KOAP, ILPO measures differences in the observation space without linear latent states and focuses on discrete actions; its training loss requires marginalization over latent actions. Similar to KOAP, LAPO jointly models forward and inverse dynamics to learn latent actions, but it regularizes latent action learning using a vector-quantized information bottleneck[23]. As vector quantization has difficulty scaling to high-dimensional continuous actions[24], KOAP takes a different approach by lifting the nonlinear system into a linear space, where the latent actions can be induced to directly capture state differences.
III Preliminaries
III-A Problem Formulation
We aim to learn a policy $\pi$ given a set of action-free expert demonstrations: $\mathcal{D}_{x}=\{(x^{i}_{0},x^{i}_{1},\cdots,x^{i}_{T})\}_{i=1}^{N}$, alongside a smaller set of demonstrations with action labels: $\mathcal{D}_{a}=\{(x^{i}_{0},a^{i}_{0},x^{i}_{1},a^{i}_{1},\cdots,a^{i}_{T-1},x^{i}_{T})\}_{i=1}^{M}$, where $M\ll N$. We wish to exploit the large action-free dataset to facilitate learning when only a small number of action-labeled trajectories are available.
III-B Plan-then-Control with Diffusion Planners and Inverse Dynamics Controllers
We adopt a plan-then-control scheme, where a planner generates $k$ future states conditioned on the current state $x_{t}$ and history $h_{t}$ (e.g., previous states): $\pi_{PL}(x_{t+1:t+k}|x_{t},h_{t})$. The inverse dynamics controller predicts actions that would give rise to these generated future states: $\pi_{C}(a_{t:t+k-1}|x_{t:t+k},h_{t})$. The combined system forms a policy $\pi(a_{t:t+k-1}|x_{t},h_{t})$ that generates action sequences conditioned on current state and history.
Our work makes use of diffusion model planners, which have been shown to be effective for planning[1, 25]. Early works have used diffusion models to directly predict actions [26, 27, 1, 28, 25, 29], but here, we diffuse intended future states, similar to [1]. This enables us to train the planner using only the action-free dataset $\mathcal{D}_{x}$. The remaining challenge and main problem addressed in this work is to learn an effective controller $\pi_{C}$ given limited action labels.
III-C Koopman Operator Theory
In our work, we learn representations with linear dynamics, which finds connections to Koopman Operator Theory[30, 31]. The key notion is to represent a non-linear dynamical system as a linear system by projecting (lifting) the states into an infinite-dimensional space.
More formally, let $x_{t}\in\mathcal{X}\subset\mathbb{R}^{n}$ denotes the state of a dynamical system at time $t$. Given a discrete-time dynamical system governed by $x_{t+1}=\mathbf{F}(x_{t})$, the Koopman Operator $\mathcal{K}:\mathcal{F}\rightarrow\mathcal{F}$ is defined as a linear transformation acting on observable functions $g$, where $\mathcal{K}g\triangleq g\circ\mathbf{F}$. Here $g:\mathcal{X}\rightarrow\mathbb{R}\in\mathcal{F}$ is the collection of all functions (also called observables) that forms an infinite-dimensional Hilbert Space. For any observable function $g$, there exists a Koopman Operator $\mathcal{K}$ such that,
$\mathcal{K}g(x_{t})=g(\mathbf{F}(x_{t}))=g(x_{t+1}).$ (1)
Koopman Operators have been extended to nonlinear control systems $x_{t+1}=\mathbf{F}(x_{t},a_{t})$ by linearly injecting the control input $a_{t}$ into the Koopman Observation space  [32, 33],
$g(x_{t+1})=\mathcal{K}g(x_{t})+\mathcal{C}a_{t}$ (2)
where $\mathcal{C}$ is referred as the Control Matrix.
In practice, by identifying a finite-dimensional subspace where $\mathcal{K}$ remains invariant, the Koopman Operator can be represented as a Koopman Matrix with finite dimension, $\mathcal{K}\in\mathbb{R}^{m\times m}$, where $g:\mathcal{X}\rightarrow\mathbb{R}^{m}$ can be hand-specified or learned[2, 34]. Deep Koopman Operators based on neural networks have been popular and applied successfully to various problems in robotics and control, including multi-object systems and soft robots[35], long-horizon planning [36], visual-motor policy learning [37], dexterous hand manipulation [38], optimal control [39], and inverse optimal control [40]. Existing learning methods typically assume the availability of abundant state-action data to learn the Koopman Operators and control matrices jointly. Next, we will extend this framework to scenarios where action data is limited.
IV Imitation Learning with Limited Action Data
In this work, we present an efficient controller assuming abundant state observations and limited action data. Our approach is to learn useful latent actions that can be quickly mapped to continuous real robot actions.
KOAP embodies a simple strategy: regularizing latent action learning with linear latent dynamics and linear action decoding. Enforcing linear forward dynamics in a lifted latent state space helps the learned latent actions capture state differences. This is in contrast to prior works like ILPO and LAPO, which model more complex relationships between latent actions and dynamics. Further regularizing the action decoder (e.g., to be linear) minimizes the risk of overfitting, particularly with limited action-labeled data.
Lifting the System and Latent Action Prediction. We use Deep Koopman embeddings to lift the system and assume that sequences of consecutive latent states encode action/control information. Let $x^{\tau}_{t}$ represent the trajectory around time $t$, consisting of past, current, and future states: $x^{\tau}_{t}=(x_{t-n},x_{t-n+1},\cdots,x_{t},x_{t+1},\cdots,x_{t+k})$. The state trajectory encoder (latent action predictor) $f_{\theta}$ maps these trajectories to control signals in the Koopman observation space:
$g_{\theta}(x_{t+1})$ $=\mathcal{K}_{\theta}g_{\theta}(x_{t})+f_{\theta}(x^{\tau}_{t})$ (3)
This linear structure encourages $f_{\theta}$ to learn the difference between lifted states, $g_{\theta}(x_{t+1})$ and $\mathcal{K}_{\theta}g_{\theta}(x_{t})$, without needing complex information bottlenecks.
We assume state-action trajectories are generated by the same system, so the observation function $g_{\theta}$ remains the same. We define an action decoder $d_{\phi}$:
$a_{t}=d_{\phi}(f_{\theta}(x^{\tau}_{t}))$ (4)
A linear decoder helps avoid overfitting when few action-labeled trajectories are available. While a natural question arises regarding the sufficiency of a linear model to remap latent actions to real robot actions, our experiments demonstrates linear decoding to be effective. However, more complex decoders can be introduced if performance declines (with the corresponding risk of overfitting).
Training. We jointly train the Koopman observation function $g_{\theta}$, Koopman matrix $\mathcal{K}_{\theta}$, and latent action predictor $f_{\theta}$. If action data is available, we also train the action decoder $d_{\phi}$. The observation function $g_{\theta}$ is trained as part of an autoencoder, with $g_{\theta}$ as the encoder and $g_{\psi}$ as the decoder:
$\mathcal{L}_{recon}=\frac{1}{T}\sum_{t=1}^{T}\|g_{\psi}(g_{\theta}(x_{t}))-x_{t}\|^{2}$ (5)
We use a Koopman dynamics loss to learn $\mathcal{K}_{\theta}$ and trajectory encoder $f_{\theta}$,
$\mathcal{L}_{kpm}=\frac{1}{T}\sum_{t=1}^{T}\|\mathcal{K}_{\theta}z_{x_{t}}+f_{\theta}(x_{t}^{\tau})-g_{\theta}{x_{t+1}}\|^{2}$ (6)
The action decoder $d_{\phi}$ is trained with an action prediction loss:
$\mathcal{L}_{a}=\frac{1}{T}\sum_{t=1}^{T}\|d_{\phi}(f_{\theta}(x_{t}^{\tau}))-a_{t}\|^{2}$ (7)
The total loss combines these terms: $\mathcal{L}_{total}=\mathcal{L}_{recon}+\lambda_{1}\mathcal{L}_{kpm}+\lambda_{2}\mathcal{L}_{a}$, where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters. If no action labels are available, the action prediction loss $\mathcal{L}_{a}$ is omitted. We apply L2 regularization to all parameters.
V Experiments
In this section, we report our experiment results using simulation benchmarks. We are interested in KOAP’s effectiveness under varying amounts of action data and hypothesize that KOAP effectively learns latent actions, which results in good policy performance when action-labeled trajectories are limited.
V-A Benchmark Tasks
We use the Datasets with Diverse human Demonstrations for Imitation Learning (D3IL) benchmark dataset [5], which contains multi-modal human demonstrations and robot trajectories for various manipulation tasks. We used five tasks of varying difficulty:
  • Avoiding: The robot’s task is to move its end-effector from left to right across a table while avoiding fixed obstacles. The robot’s state is defined by the position of its end-effector, and its actions are based on controlling its end-effector velocity. Any collision with an obstacle leads to immediate failure.
  • Aligning: In this task, the robot must manipulate a block to match a specified target pose. The state consists of the robot’s end-effector position, the block’s pose, and the target pose, while actions control the end-effector’s velocity. The goal is to push or pull the block until it reaches the target configuration.
  • Sorting (4 blocks): The robot must sort four blocks (two red, two blue) by pushing them into their designated zones. The state includes the robot’s end-effector position and the pose of the blocks, with actions controlling the end-effector’s velocity. Failure occurs if a block is pushed into the wrong zone.
  • Stacking: This is a challenging task, requiring the robot to stack blocks in a specific order within a target zone. The state includes the end-effector position, gripper status (open or closed), block poses, and the required stacking order. Actions involve controlling the robot’s 7-DoF joint velocity (which is not observed) and the gripper. Placing blocks outside the target zone or in the wrong order results in failure.
  • Partially-Observable Sorting: A more difficult partially-observable version of original Sorting task, where the exact poses of the blocks are hidden and only their positions can be observed. In addition, Gaussian noise is added to robot’s end-effector positions in the state.
Due to variation in dataset size across tasks, for training, we use 50% of the action-labeled data for the smaller task datasets (avoiding and aligning) and 25% for larger datasets (sorting and stacking). We defined five different “levels” of action data based on dataset size. For the smaller tasks, the levels corresponded to (2%, 5%, 10%, 25%, 50%), while for the larger tasks, the levels were (1%, 2%, 5%, 10%, 25%).
Figure 2: Original D3IL Tasks. The robot learns from a dataset containing partially-labeled multi-modal expert demonstrations. The goal is to manipulate object(s) to reach target positions or poses, adhering to task-specific rules (e.g., collision-free or color-based sorting).
V-B Baselines
In addition to KOAP, we ran several strong baseline methods on the above tasks:
  • Diffusion Policy (DP) [27]: A popular diffusion-based method for learning multi-modal behaviors, which uses action-labeled trajectories for training.
  • Decision Diffuser (DD) [1]: This method is similar to KOAP, but uses a pure supervised-learning approach for learning the controller. It allows us to better isolate the benefit of improving the inverse dynamics controller in a decision diffuser.
  • VAE [41]: A Variational Autoencoder method that trains a controller by encoding observational trajectories into an embedding. After training the VAE, an action decoder is trained to map sampled latent embeddings from the VAE to robot actions. This model can potentially learn useful action embeddings from the observations, but does not leverage dynamics.
  • LAPO [4]: We adopt the Decision Diffuser as above, but use LAPO to learn the inverse dynamics controller. Since no official implementation of LAPO is publicly available, we implemented LAPO following the paper with one key difference: the original LAPO paper uses VQ-VAE [42], but we instead employ the more recent Finite Scalar Quantization (FSQ) method [24], which scales better to larger codebooks and is more suitable for continuous action spaces; in preliminary trials, FSQ outperformed VQ-VAE. All observational data is used for learning the dynamics models, while action-labeled data is used to learn the mapping from latent actions to real robot actions.
All methods, except for Diffusion Policy, follow a plan-then-control scheme. All plan-then-control methods use the same Decision Diffuser for a given task. For fair comparison, we use the same underlying model for the Diffusion Policy and Decision Diffuser, with the difference being that the Diffusion Policy predicts future actions instead of states. For KOAP, $g_{\theta}$ and $g_{\psi}$ were MLPs and $f_{\theta}$ was a LSTM. The decoder was $d_{\phi}$ was a linear model.
We use the original dataset provided by D3IL and sample fixed-length trajectory sequences. The first two states (no actions) are used as conditioning input for the planner and controller. The controllers are trained to predict a sequence of actions (12 steps). During inference, we employ receding-horizon control: the first four predicted actions are executed in the environment, with replanning and execution until the episode is completed. For complete implementation details, please refer to our online code repository [^1] To be made available after review..
Figure 3: Average the success rates of each method versus the relative amount of action data used.
TABLE I: Task success rates for the different methods given varying amounts of action data (Levels 1 to 5 from smallest to largest). For Stacking, we report success rate of stacking 1 block at target position.
Action Data Amt. Task DP DD VAE LAPO KOAP (ours)
1 Avoiding $6.93\pm 4.55$ $36.70\pm 11.80$ $36.95\pm 15.84$ $43.63\pm 12.57$ $\mathbf{78.03\pm 8.63}$
Aligning $10.55\pm 1.08$ $23.90\pm 10.92$ $38.00\pm 4.33$ $36.97\pm 5.15$ $\mathbf{80.28\pm 4.94}$
Sorting $0\pm 0$ $0\pm 0$ $36.80\pm 8.03$ $11.38\pm 14.81$ $\mathbf{61.65\pm 5.63}$
Stacking $12.22\pm 3.42$ $12.22\pm 9.70$ $7.78\pm 5.82$ $6.67\pm 5.36$ $\mathbf{27.22\pm 8.89}$
PO-Sorting $0\pm 0$ $0\pm 0$ $8.62\pm 4.36$ $1.11\pm 2.48$ $\mathbf{45.83\pm 6.85}$
2 Avoiding $10.82\pm 2.10$ $77.78\pm 7.11$ $68.05\pm 14.63$ $70.82\pm 12.66$ $\mathbf{98.05\pm 1.13}$
Aligning $45.00\pm 4.61$ $72.23\pm 7.93$ $67.72\pm 9.36$ $60.85\pm 8.75$ $\mathbf{85.55\pm 2.47}$
Sorting $0.28\pm 0.63$ $49.67\pm 7.57$ $43.28\pm 7.66$ $25.55\pm 16.79$ $\mathbf{67.23\pm 7.03}$
Stacking $24.18\pm 4.00$ $21.4\pm 9.90$ $9.17\pm 3.68$ $14.18\pm 9.54$ $\mathbf{26.38\pm 13.53}$
PO-Sorting $0\pm 0$ $1.67\pm 2.35$ $13.88\pm 3.14$ $6.11\pm 5.50$ $\mathbf{48.89\pm 5.67}$
3 Avoiding $21.12\pm 6.44$ $70.80\pm 9.07$ $64.18\pm 4.27$ $76.38\pm 6.41$ $\mathbf{97.22\pm 2.49}$
Aligning $63.33\pm 8.27$ $83.62\pm 4.00$ $68.00\pm 8.81$ $75.57\pm 7.37$ $\mathbf{86.12\pm 3.14}$
Sorting $5.28\pm 2.23$ $\mathbf{64.38\pm 5.05}$ $55.80\pm 3.30$ $33.07\pm 18.90$ $61.38\pm 6.07$
Stacking $28.60\pm 5.31$ $30.55\pm 7.32$ $16.38\pm 7.23$ $20.57\pm 12.06$ $\mathbf{33.88\pm 7.73}$
PO-Sorting $2.78\pm 2.26$ $27.50\pm 8.94$ $16.38\pm 1.93$ $30.55\pm 4.04$ $\mathbf{51.95\pm 4.35}$
4 Avoiding $78.90\pm 7.68$ $88.05\pm 4.45$ $77.77\pm 9.68$ $75.00\pm 4.59$ $\mathbf{98.33\pm 1.92}$
Aligning $77.77\pm 7.98$ $81.93\pm 2.80$ $83.80\pm 4.11$ $82.50\pm 5.51$ $\mathbf{88.35\pm 3.19}$
Sorting $21.67\pm 6.08$ $62.72\pm 6.57$ $61.23\pm 2.95$ $44.18\pm 15.49$ $\mathbf{68.62\pm 8.03}$
Stacking $\mathbf{42.50\pm 8.10}$ $30.55\pm 4.14$ $23.33\pm 5.36$ $27.78\pm 13.96$ $36.95\pm 10.80$
PO-Sorting $19.72\pm 5.64$ $31.67\pm 12.47$ $20.82\pm 5.59$ $31.11\pm 7.31$ $\mathbf{54.44\pm 7.80}$
5 Avoiding $94.70\pm 1.78$ $91.40\pm 4.24$ $86.37\pm 3.52$ $81.93\pm 5.05$ $\mathbf{97.50\pm 1.58}$
Aligning $81.68\pm 3.74$ $81.08\pm 7.05$ $80.23\pm 7.05$ $82.77\pm 4.70$ $\mathbf{91.95\pm 5.38}$
Sorting $42.77\pm 5.25$ $66.08\pm 4.46$ $\mathbf{67.88\pm 4.15}$ $64.17\pm 8.31$ $65.83\pm 3.68$
Stacking $\mathbf{53.05\pm 10.47}$ $32.78\pm 9.31$ $30.83\pm 9.98$ $27.20\pm 7.62$ $45.57\pm 5.43$
PO-Sorting $45.58\pm 3.82$ $45.55\pm 3.81$ $22.76\pm 2.85$ $31.11\pm 7.80$ $\mathbf{58.33\pm 5.00}$
V-C Results and Analysis
KOAP learns latent actions that are informative and results in strong policy performance. Fig. 3 and Table I summarizes each method’s policy success rate averaged across six different seeds, with 60 rollouts per seed (a total of 360 rollouts). KOAP significantly outperforms the standard Decision Diffuser, highlighting the value of learning latent actions from observations. Performance differences were most pronounced when action data was scarce ($<$ 5%), with the gap narrowing as more action data became available, which is consistent with the expectation that more data allows for better inverse model learning. In PO-Sorting, the baselines experience significant degradation relative to their performance in Sorting, while KOAP retains good performance. This indicates that KOAP learns meaningful latent actions, even under partially-observability.
For these D3IL tasks, LAPO did not perform well, especially on the high-dimensional Stacking task, potentially because information valuable for precise continuous action control was lost in the quantization bottleneck. In contrast, KOAP uses continuous representations throughout, with regularization imposed in the linear latent dynamics. Moreover, the VAE learner’s relatively lower performance suggests that incorporating dynamics information is important for learning good latent actions.
KOAP learns better latent actions given more observational data. Figure 4 demonstrates a clear increase in success rates as more observation data is provided, while holding the action labels constant at 1%. This result suggests that sufficient observation data is necessary for effective lifting of the nonlinear system and obtaining quality representations for latent action learning.
Figure 4: KOAP’s performance increases with observation data.
Comparison with Variants. To better understand how KOAP’s design influences its performance, we conducted experiments with several variants:
  • Nonlinear: Similar to KOAP but with nonlinear latent dynamics, where $g(x_{t+1})=T_{\theta}(g(x_{t}),f(x^{\tau}))$, and $T_{\theta}$ is a neural network. It also uses a noninear action decoder.
  • Pretrain: A two-stage approach where KOAP’s latent actions are first trained without any action labels. Then, action labels are used to train the decoder and finetune KOAP’s latent models (without access to the original observation data).
  • Relabel: Inspired by ILPO [22], this variant learns an action predictor without a forward dynamics model and uses it to relabel the dataset, which is then used to train a standard Decision Diffuser.
TABLE II: Comparison of KOAP Variants showing Average Success Rates
Action Data Amt. Task KOAP Nonlinear Pretrain Relabel
1 Avoiding $78.03\pm 8.63$ $64.17\pm 13.97$ $82.22\pm 7.44$ $5.83\pm 10.96$
Stacking $27.22\pm 8.89$ $14.17\pm 8.86$ $28.33\pm 10.27$ $4.17\pm 6.07$
2 Avoiding $98.05\pm 1.13$ $70.00\pm 10.80$ $94.17\pm 6.07$ $57.50\pm 11.09$
Stacking $26.38\pm 13.53$ $21.67\pm 4.71$ $27.50\pm 7.50$ $7.50\pm 5.59$
3 Avoiding $97.22\pm 2.49$ $83.33\pm 8.50$ $98.33\pm 2.36$ $65.83\pm 16.18$
Stacking $33.88\pm 7.73$ $26.67\pm 8.98$ $32.50\pm 9.47$ $16.67\pm 9.86$
4 Avoiding $98.33\pm 1.92$ $96.67\pm 3.73$ $98.33\pm 2.36$ $71.67\pm 12.13$
Stacking $36.95\pm 10.80$ $38.33\pm 4.71$ $38.33\pm 10.27$ $19.17\pm 14.55$
5 Avoiding $97.50\pm 1.58$ $99.17\pm 1.86$ $93.33\pm 12.80$ $75.83\pm 6.07$
Stacking $45.57\pm 5.43$ $39.17\pm 13.36$ $47.50\pm 9.46$ $19.17\pm 5.34$
Table II shows success rates (averaged over 6 seeds, each with 20 rollouts per method/task) for these variants on two D3IL tasks. We observe that employing a nonlinear latent dynamics model was not as effective, potentially because more parameters were introduced and the latent action learner was not longer forced to learn differences between latent states. The relabeling variant also underperformed, reinforcing the idea that observational data enhances latent action learning. Notably, even without any action labels, KOAP was able to learn useful prior models that were easily finetuned when action data becomes available.
VI Case Study: Real Robot Scooping
We evaluate KOAP on a robot scooping task, where the goal is to use a spoon to manipulate a chocolate piece in a bowl. The experiment was conducted using a 7-DoF Panda robot.
Human demonstrations were collected by manually scooping the chocolate with the spoon, and robot action-labeled demonstrations were gathered through kinesthetic teaching using the same spoon. We collected 100 trajectories for both human and robot demonstrations. To reduce the computational overhead of generating pixel states for planning, we use a pretrained ResNet[43] to extract feature embeddings from pixel observations. The diffusion planner generates future embeddings conditioned on the current pixel observations, and the controller produces robot joint angles as actions.
Figure 5: Real robot experiment setup: The robot starts by holding a spoon in a random initial position near the bowl. A camera provides a third-person view of the scene. Using pixel observations, the robot attempts to scoop the chocolate into the target container. We use two evaluation metrics: rim, which indicates the robot was able to push the chocolate to the rim of the bowl, and success, which indicates the chocolate was successfully scooped into the target container.
Figure 6: Performance across three different action data settings for the scooping task. We report success rates for both pushing the chocolate to the rim of the bowl and successfully scooping it into the target container.
For each configuration, we ran 15 test episodes. Three random initial spoon positions were selected, and five episodes were conducted for each starting position. The success rates for the scooping task are shown in Fig. 6.
The results show that KOAP outperforms the baseline methods in both rim and success rates across varying amounts of action-labeled trajectory data. It significantly improves the standard Decision Diffuser and LAPO by learning a better controller. We observed the Diffusion Policy performed well at specific initial spoon configurations, but was not able to generalize to other locations. While there is room for improvement for all methods, these results shows that KOAP is able to leverage action-free demonstrations to enhance real robot performance.
VII Conclusion
In this work, we proposed the Deep Koopman Operator with Action Proxy (KOAP) method to enhance action-data efficiency in policy learning by leveraging observational data. Through simulated robot manipulation tasks, we demonstrated that KOAP effectively utilizes observational data to improve action-data efficiency in policy learning. In addition, our real-world robot experiments showed that KOAP can successfully incorporate visual human demonstrations to facilitate visual-motor policy learning with minimal robot kinesthetic teaching trajectories. KOAP opens a new perspective for data-efficient policy learning, which could significantly contribute to the development of general-purpose robots in the future.